# SGLDBench: A Benchmark Suite for Stress-Guided Lightweight 3D Designs

### v1

@@ Abstract
We introduce the Stress-Guided Lightweight Design Benchmark (SGLDBench), a comprehensive benchmark suite to apply and evaluate material layout strategies for generating stiff lightweight designs in 3D domains. SGLDBench provides a seamlessly integrated simulation and analysis framework, providing six reference strategies accompanied by a scalable multigrid elasticity solver to efficiently execute these strategies and validate the stiffness of their results. This facilitates systematic analysis and comparison of design strategies regarding the mechanical properties they achieve. SGLDBench enables the evaluation of diverse settings of load conditions and, through the tight integration of the solver, enables support for high-resolution designs and stiffness analysis. Moreover, SGLDBench emphasizes visual analysis to explore relations between the geometric structure of a design and the distribution of stresses, providing insights into the specific properties and behaviors of different design strategies. SGLDBenchs’ specific features are highlighted in several experiments, by comparing the results of reference strategies with respect to geometric and mechanical properties.
@@ 1 Introduction
Topology optimization and functionally graded lattice infill are primary strategies for designing mechanically sound lightweight structures, i.e., structures with high stiffness when loaded via external forces. Topology optimization [1, 2] transforms the design process into an optimization problem of determining where to place material. Functionally graded lattice infill can be considered a rational distribution of lightweight (meta-)material with a geometric representation [3].
Topology optimization usually does not consider the geometric structure of the resulting material layout. Lattice infill, on the other hand, uses a structure made of individual edges (the struts), forming a graph or a grid composed of polyhedral cells (the scaffold). For designing load-bearing structures, tailoring the layout of lattice infill based on the stress distribution is key to achieving high structural performance. In particular, the structural rigidity of an infill is increased if the material is aligning with orthogonal principal stress directions in the object under load [4]. In the limit of material volume, this alignment forms micro-structures resembling quads or hexahedra [5].
Figure 1: Motivational example. (a) Problem description. The greyish region shows the design domain. Black rings indicate fully solid passive elements. A constant load exerts on all inner boundary vertices of the upper ring (green). All inner boundary vertices of the bottom ring (orange) are fixed. (b) density-based topology optimization [6]. (c) Porous infill optimization [7]. (d) Voronoi infill [8]. (e) Stress-line-guided material layout [9]. (f) Conforming lattice structures [10]. (g) Volumetric Michell trusses  [11]. All designs consume roughly the same amount of material. $c$ refers to the compliance of the design, which is measured by the strain energy.
@@
Many different strategies for generating lightweight structures have been proposed over the last years [12, 3, 13, 14], the majority of them demonstrated on 2D domains. While the extension of topology optimization to 3D is rather straightforward, the generation of 3D-domain-filling lattice structures that follow the major stress directions is challenging. This is in particular due to the existence of degenerate points[15] (or degenerate regions in 3D domains  [16, 17, 18]) where the stress tensor has repeating eigenvalues and the principal stress directions cannot be decided. Due to this, integrability conditions are violated and consistent domain parameterizations cannot be computed [19].
Topology optimization, in its basic form, aims at achieving the highest structural stiffness. In principle, alternative strategies share this goal, yet each has additional considerations that may come with a compromise in stiffness, such as achieving geometric properties like regularity (i.e., element type variation), uniformity (i.e., element size variation) or space-fillingness for improved robustness, as well as purely aesthetic features. To support users in selecting the best strategy for different use cases and researchers in identifying open research questions, a benchmark suite for generating, analyzing, and comparing the results of different strategies is required.
A few benchmark papers have addressed issues like special solvers for topology optimization [20], benchmarks for 2D topology optimization in specific load cases [21], practices that should be considered when performing topology optimization [22], as well as the mechanical soundness of simple lattice infills like orthogonal grids and shells [23]. Different architectures for multidisciplinary design optimization have been reviewed and compared [24], and issues of design and structural optimization of lightweight design have been discussed, especially in the context of additive manufacturing [12]. A review of uniform and non-uniform lattice structures such as foams and honeycombs sheds light on their properties and methods for designing and optimizing such structures [25, 26]. Unit cell lattices comprising structures made of a single type of cells have been researched [27], and the properties of certain types of lattice infills regarding 3D printing processes have been discussed [28]. The combination of topology optimization and micro element-based lattice infills have resulted in structures exhibiting anisotropic mechanical properties[29]. A benchmark enabling researchers and users to efficiently compute 3D designs with different strategies and effectively compare the results regarding their mechanical and structural properties does not exist.
We introduce SGLDBench to close this gap. It comprehensively investigates the combination of boundary shapes and conditions with lightweight design strategies in 3D domains. We demonstrate SGLDBench with reference strategies, including comparisons regarding mechanical and geometric properties. This is made possible by a seamlessly integrated, MATLAB-based simulation framework, which implements a number of reference strategies. To facilitate an effective comparison of different designs, SGLDBench generates a scalar material field regardless of the used strategy, and provides options to analyze the stiffness and structure of this field.
SGLDBench offers the following key features:
  • Selection of Lightweight 3D Design Strategies: The benchmark offers different strategies that allow for drawing analogies between topology optimization and lattice infill, and enable comparative studies with new scenarios and designs.
  • Material Layout Generation: Central to the benchmark is the voxelization of complex infills into a Cartesian simulation grid, to enable a consistent comparison of different lightweight designs—given as a material field, a mesh, or just an edge-graph—regarding their mechanical properties.
  • Simulation Suite: SGLDBench provides a MATLAB-interfaced simulation framework including an efficient multigrid solver to generate high-resolution stress fields and efficiently assess the stiffness of a design.
  • Visual Design Analysis: The benchmark is accompanied by a fast volume visualization module, providing visual feedback regarding a design’s shape, and the stress directions and material alignment before and after layout optimization.
We chose MATLAB as working environment due to its widespread use in computational design. SGLDBench utilizes MATLAB’s simulation and visualization operations for design generation and analysis. New design strategies can be integrated, either by MATLAB programs or by using MATLAB’s functionality to call executables or Python programs from other code bases via inline calls. All SGLDBench specific operations have been either implemented in MATLAB or are publicly available MATLAB or C++/Python programs. SGLDBench calls publicly available external libraries for core operations such as meshing and voxelization. The visualization module has been implemented in WebGL to enable its use either as a standalone viewer in a web-browser, or via inline calls to MATLAB’s viewing functionality. Upon acceptance, the entire code basis of SGLDBench will be made publically available.
2 From Boundary Conditions to Material Layouts
We first discuss a motivational study in 2D that illustrates several key aspects of SGLDBench. The user specifies the boundary conditions, i.e., the boundary of a domain as well as the exerting loads and fixed boundary parts, see Fig. 1a. The domain is then voxelized, i.e., it is discretized with a Cartesian grid, and per-voxel properties are set according to the boundary conditions and material properties. We call this configuration a preset.
With the selected preset, SGLDBench simulates the object’s internal stress field using a multigrid finite-element elasticity solver [30, 31, 32, 33]. While some design strategies require repeated solver executions to iteratively optimize the material layout, some others generate an infill structure that is guided by the principal stresses in the initial field.
The user can select from six layout strategies for computing an infill design: topology optimization [34], porous infill optimization [7], Voronoi infill [8], stress-line-guided material layout [9], conforming lattice structures [10] and volumetric Michell trusses [11].
@@ 2 From Boundary Conditions to Material Layouts
Figure 2: Design strategies producing a 3D material field: (a) Density-based topology optimization [6]. (b) Porous infill optimization [7]. Design strategies producing a graph structure: (c) Voronoi infill [8]. (d) Stress-line-guided material layout [9]. (e) Conforming lattice structures [10]. (f) Volumetric Michell trusses  [11]. All designs consume roughly the same volume fraction of 0.4, and they are subject to the same boundary conditions, i.e., the bottom of the design domain is fixed, and the loads are indicated by green arrows in (a). Below each design is its compliance. SGLDBench uses isosurface volume rendering with ambient occlusion for visualization. All designs are coated with fully solid boundary elements with a consistent thickness. In (c) to (f), these elements are peeled away to show the infills.
@@
Our selection has not been made with the intention to favor any of these methods, but to reveal the specific characteristics of 3D design strategies following different objectives. The methods span a wide spectrum from purely compliance-based optimization to geometry-aware infill generation. We select density-based topology optimization as representative of various topology optimization approaches. It serves as a reference for the compliance that can be achieved. Porous infill optimization, while not explicitly using the principal stress directions, results in wall-like structures that largely agree with two of these directions. Voronoi infill considers only the stress magnitude but not its principal directions. In contrast, material layouts guided by the Principal Stress Lines (PSLs) follow exactly the principal stress trajectories in the initial solid domain. PSL-guided infills serve as a reference conveying these directions, even though the final designs are not connected in general. Conforming lattice structures and volumetric Michell trusses aim at finding a balance between stress alignment and geometric regularity of the resulting designs. While the first approach favors stress alignment and, therefore, needs to resort to an edge-graph structure, the latter approach strives for a pure hexahedral mesh and, therefore, needs to sacrifice stress alignment. Figs. 1b-g show the infill designs computed by these methods with our motivational 2D example.
@@
While topology optimization and porous infill generate a material field, all other methods compute a lattice structure composed of edges and nodes. SGLDBench can show lattice structures via MATLAB’s visualization functionality, by rendering edges as cylinders with a selected width. In addition, SGLDBench voxelizes each optimized lattice structure into a material field on a Cartesian grid with the same resolution as the initial preset. For a final material field and boundary conditions, the linear elasticity solver computes the compliance of the design, i.e., the inverse stiffness or strain energy. Notably, the user can also upload a material field and define a preset to perform the analysis with SGLDBench. Thus, the user can assess the robustness of a design with respect to loads it has not been optimized for.
SGLDBench supports different visual analysis options to inspect a 3D design. The geometry of a design is visualized via isosurface rendering of the 3D material field—in combination with a view-plane-aligned clip-plane—to convey a design’s uniformity and regularity. For a 3D human femur under load, Fig. 2 shows visualizations of the infill designs that are produced with SGLDBench’s methods, together with their compliances when loaded with the specific boundary forces. The compliance of each design reveals the relationship between stiffness and structure. The 3D stress fields produced by SGLDBench can be visualized with trajectory-based visualization techniques available in MATLAB [35]. Over that, SGLDBench provides tailored visualization options to investigate how the mechanical properties of the initial solid and the generated infill design change. Direct volume rendering is used with predefined color and opacity transfer functions to reveal the per-voxel deviation of the stresses in a final design from the stresses in the initial solid body (see Fig. 3).
Figure 3: Exemplary visual analysis of stresses in (a) a PSL-guided infill and (b) a Voronoi infill. Per voxel deviation between the major stress direction under the same load in the initial solid object and the infills is color-coded from light brown (low) to dark brown (high). The PSL-guided infill follows the initial stress directions and produces similarly directed stresses. Voronoi infill edges in the object interior are not related to the initial stress directions. Edges going through the passive boundary elements (which are the same in the solid and the voxelized infill) have approximately the same stresses.
@@ 3 Components of SGLDBench
We describe here the most important features and operations of SGLDBench, including descriptions of the supported topology optimization and lattice infill methods. The use of SGLDBench is demonstrated in the accompanying video.
@@ 3.1 Domain specification
Voxelization. The simulation grid is computed by voxelizing the simulation domain. The user inputs the domain boundary as a closed triangular mesh. SGLDBench uses MATLAB’s voxelization capabilities  [36] with a user-defined voxel resolution to compute a solid voxelization. Alternatively, the user can upload a 3D voxel grid discretizing the domain and flagging each voxel representing the object as solid or void otherwise. All voxels with at least one exterior voxel in their 26-neighborhood are considered boundary voxels. SGLDBench provides a dilation operation to thicken the current boundary, setting each exterior voxel with at least one boundary voxel in its 26-neighborhood as boundary. The voxelized object is the basis of all subsequent operations in SGLDBench.
Boundary Conditions. To conduct a linear elasticity simulation of an object using FEA, the boundary conditions need to be specified, i.e., where the object is fixed, and where and how the object is loaded. Fixations and forces are applied at nodes of the boundary elements. The user defines the extent and position of an axis-aligned box (or a sphere), and SGLDBench automatically fixes or assigns a specified load to all boundary nodes within this object. In the same way, the nodes which should be reset can be selected. When all nodes of a finite element are fixed, the entire element is fixed, i.e., it becomes rigid and does not respond to any loads.
Passive Elements. In many practical design and optimization scenarios, passive elements are used to maintain certain geometric features, like the object boundary or notches in a machine part where connections will be mounted. A passive element is constantly solid during design and optimization, and it contributes to the stiffness of the design. SGLDBench considers two general situations for specifying the passive elements. One is to set all boundary elements to passive, corresponding to a common infill design problem. Another one is to set the elements with loaded or fixed nodes to passive, to maintain certain prescribed features after design and optimization. By using a dilation operation, passive structures can be enlarged. Fig. 4 shows the transformation from a boundary mesh to a voxel model with boundary conditions and passive elements.
Figure 4: From the design domain to the finite element model with boundary conditions. (a) A triangle mesh defines the domain boundary. (b) After voxelizing the domain, the user specifies boundary forces and fixations at grid vertices. Black and violet dots, respectively, indicate fixed vertices and vertices loaded with a constant downward force. (c) Passive elements are shown in brown.

### v2

@@ Abstract
We introduce the Stress-Guided Lightweight Design Benchmark (SGLDBench), a comprehensive benchmark suite for applying and evaluating material layout strategies to generate stiff, lightweight designs in 3D domains. SGLDBench provides a seamlessly integrated simulation and analysis framework, including six reference strategies and a scalable multigrid elasticity solver to efficiently execute these strategies and validate the stiffness of their results. This facilitates the systematic analysis and comparison of design strategies based on the mechanical properties they achieve. SGLDBench enables the evaluation of diverse load conditions and, through the tight integration of the solver, supports high-resolution designs and stiffness analysis. Additionally, SGLDBench emphasizes visual analysis to explore the relationship between the geometric structure of a design and the distribution of stresses, offering insights into the specific properties and behaviors of different design strategies. SGLDBench’s specific features are highlighted through several experiments, comparing the results of reference strategies with respect to geometric and mechanical properties.
@@ 1 Introduction
Topology optimization (TO) and functionally graded lattice infill are primary strategies for designing mechanically sound, lightweight structures, i.e., structures with high stiffness (corresponding to a low compliance, or degree of deformability) under applied loads. TO determines the optimal material distribution within a given design domain to achieve a desired structural performance, such as maximizing stiffness, while satisfying constraints like material use [1, 2]. Functionally graded lattice infill refers to a design approach in which the density, size, shape, or material properties of the lattice structure vary spatially within a 3D object to meet specific performance requirements [3]. For beam-based lattices, the final design can be represented by a graph or grid composed of polyhedral cells, each constructed from individual edges.
TO, in its basic form, does not consider the geometric structure of the resulting material layout but aims to achieve the highest possible stiffness. Lattice infill design strategies, in principle, share this goal, by tailoring the lattice layout based on the stress distribution. Stress is a measure of the internal forces that develop within a material when it is subjected to external loads and quantifies the intensity of these forces at a specific point in the material. The structural rigidity of an infill increases when the material aligns with the orthogonal principal stress directions of the object under load [4]. These directions, corresponding to the eigenvectors of the 3x3 stress tensor, indicate the normal stresses acting on specific planes within where shear stresses are zero. At the limit of material volume, considering these directions for the lattice layout results in microstructures resembling quads or hexahedra [5, 6].
Each lattice infill design strategy, however, involves additional considerations that may compromise stiffness. These include achieving geometric properties such as regularity (i.e., variation in element type), uniformity (i.e., variation in element size), or space-fillingness to enhance robustness, as well as purely aesthetic features [7, 3, 8, 9].
@@
Moreover, generating 3D domain-filling lattice structures that align with major stress directions presents significant challenges. This difficulty arises from the existence of degenerate points [10] (or degenerate regions in 3D domains [11, 12, 13]) where the stress tensor has repeating eigenvalues, making the principal stress directions indeterminate. As a result, integrability conditions are violated, and consistent domain parameterizations cannot be computed [14].
To assist users in selecting the right 3D lightweight design strategy for various use cases, and to help researchers identify open research questions, a benchmark suite for generating, analyzing, and comparing the results of different strategies is essential.
@@
Figure 1: Infill designs computed using the strategies provided by SGLDBench. All designs consume roughly the same amount of material and are subject to the same boundary conditions: the bottom of the design domain is fixed, and the loads are indicated by green arrows in (a). Below each design, its compliance $c$ is provided. For visualization, isosurface volume rendering is used. All designs are coated with fully solid boundary elements of consistent thickness. In (c) to (f), these elements are peeled away to reveal the infills.
A few benchmark papers have addressed issues such as special solvers for TO [15], benchmarks for 2D TO in specific load cases [16], practices that should be considered when performing TO [17], as well as the mechanical soundness of simple lattice infills such as orthogonal grids and shells [18]. Different architectures for multidisciplinary design optimization have been reviewed and compared [19], and the design and structural optimization of lightweight design has been discussed, especially in the context of additive manufacturing [7]. A review of uniform and non-uniform lattice structures such as foams and honeycombs sheds light on their properties and methods for designing and optimizing such structures [20, 21]. Unit cell lattices comprising structures made of a single type of cells have been researched [22], and the properties of certain types of lattice infills regarding 3D printing processes have been discussed [23]. The combination of TO and micro element-based lattice infills have resulted in structures exhibiting anisotropic mechanical properties [24, 25]. There is no benchmark that allows researchers and users to efficiently compute 3D designs with different strategies and to effectively compare the results with respect to their mechanical and structural properties.
We introduce SGLDBench to address this gap. It provides a comprehensive investigation of the combination of boundary shapes and conditions with lightweight design strategies in 3D domains. This is facilitated by a seamlessly integrated, MATLAB-based simulation and analysis framework offering the following key features:
  • Selection of Lightweight 3D Design Strategies: The benchmark includes various strategies, enabling comparisons between TO and lattice infill, as well as studies of new scenarios and designs. For a 3D human femur under load, Fig. 1 shows visualizations of the infill designs computed by SGLDBench.
  • Material Layout Generation: A central feature of the benchmark is the voxelization of complex infills into a Cartesian simulation grid. This ensures consistent comparisons of lightweight designs—whether represented as a material field, mesh, or edge graph—with respect to their mechanical properties.
  • Simulation Suite: SGLDBench provides a MATLAB-interfaced simulation framework with an efficient multigrid solver for generating high-resolution stress fields and assessing the stiffness of a design efficiently.
  • Visual Design Analysis: A fast volume visualization module accompanies the benchmark, offering visual feedback on a design’s shape, stress distribution, and material alignment before and after layout optimization.
We chose MATLAB as the working environment due to its widespread use in computational design. SGLDBench leverages MATLAB’s simulation and visualization capabilities for design generation and analysis. New design strategies can be integrated either through MATLAB programs or by using MATLAB’s functionality to call executables or Python programs from other codebases via inline calls. All SGLDBench-specific operations have been implemented in MATLAB or rely on publicly available MATLAB, C++, or Python programs. SGLDBench also uses external libraries for core operations such as meshing and voxelization. The visualization module is implemented in WebGL, allowing it to function either as a standalone viewer in a web browser or via inline calls to MATLAB’s viewing functionality. Upon acceptance, the entire codebase for SGLDBench will be made publicly available.
Figure 2: SGLDBench’s visual analysis options. (a) Major (brown) and minor (green) PSLs according to boundary conditions from Fig. 1. (b) Direct volume rendering of scalar von Mises stresses in the solid domain. (c) Voronoi infill (c = 2.86). (d) Infill with color-coded von Mises stresses. (e) Same as (d), but different forces apply and von Mises stresses change (c = 4.14). (f) Comparative visual stress analysis showing the misalignment between the major stress directions in the solid and the infill under the same loads.
2 SGLDBench’s Functional Structure
We begin by introducing the key functionality of SGLDBench. SGLDBench provides an interface to enable users to specify the boundary conditions, which include the boundary of a domain, the applied loads, and the fixed boundary regions. The domain is then voxelized, meaning it is discretized into a Cartesian grid, with per-voxel properties assigned based on the boundary conditions and material properties. We refer to this configuration as a preset.
Using the selected preset, SGLDBench simulates the object’s internal stress field via a multigrid finite-element elasticity solver [26, 27, 28, 29]. While some design strategies require repeated solver executions to iteratively optimize the material layout, others generate an infill structure guided by the principal stresses in the initial field.
Users can choose from six layout strategies for computing an infill design: density-based TO  [30], porous infill optimization [31], Voronoi infill [32], stress-line-guided material layout [33], conforming lattice structures [34], and volumetric Michell trusses [35].
@@ 2 SGLDBench’s Functional Structure
Our selection has not been made with the intention to favor any of these methods, but to reveal the specific characteristics of 3D design strategies following different objectives. The methods span the spectrum from purely stiffness-based optimization to geometry-aware infill generation. We select density-based TO as a representative of various TO approaches. It serves as a reference for the stiffness that can be achieved. Porous infill optimization, while not explicitly using the principal stress directions, results in wall-like structures that largely agree with two of these directions. Voronoi infill considers only the stress magnitude but not its principal directions. In contrast, material layouts guided by the Principal Stress Lines (PSLs) follow exactly the mutually orthogonal principal stress trajectories in the initial solid domain. PSL-guided infills serve as a reference conveying these directions, even though the final designs are not connected in general. Conforming lattice structures and volumetric Michell trusses aim at finding a balance between stress alignment and geometric regularity of the designs. While the first approach favors stress alignment and, therefore, needs to resort to an edge-graph structure, the latter approach strives for a pure hexahedral mesh and, therefore, needs to sacrifice stress alignment.
@@
While TO and porous infill generate a material field, other methods compute a lattice structure composed of edges and nodes. SGLDBench voxelizes these structures into a material field on a Cartesian grid with the same resolution as the initial preset. Using the material field and boundary conditions, the elasticity solver computes the compliance of the design. When using iterative optimization methods, the compliance history is recorded and can be visualized.
SGLDBench supports different visualization options to inspect a 3D design. The principal stress directions in a stress field are visualized using PSL-guided trajectory visualization implemented in MATLAB programs [36]. These programs are accessible through SGLDBench’s interface and allow users to customize the number and appearance of the visualized trajectories (see Fig. 2a).
From the principal stresses the scalar von Mises stresses are computed. The von Mises stress is commonly used in engineering and materials science to predict yielding in ductile materials under load. It provides a single value that reflects the combined effect of all stress components acting on a material. This scalar field can then be visualized with SGLDBench’s WebGL-based visualization module for enhanced rendering performance (Fig. 2b). An infill structure is rendered as an iso-surface in the material field (Fig. 2c), and it can be color-coded with the von Mises stress to reveal local stress concentrations (Fig. 2d).
Additionally, a variable load structural analysis can be performed, by loading a design with forces different from those for which it was initially optimized. This functionality enables users to evaluate the robustness of a design under different loading conditions. The optimized design can be color-coded with the von Mises stresses occurring under the new load conditions (Fig. 2e).
Furthermore, SGLDBench offers tailored visualization options to examine how the mechanical properties of the initial solid and the generated infill design have changed. Direct volume rendering is used with a predefined color transfer function to visualize the per-voxel stress deviations in the final design relative to those in the initial solid body (see Fig. 2f).
@@ 3 Components of SGLDBench
We describe here the most important features and operations of SGLDBench, including descriptions of the supported TO and lattice infill methods. The use of SGLDBench is demonstrated in the accompanying videos.
@@ 3.1 Domain specification
Voxelization. The simulation grid is created by voxelizing the simulation domain. The user provides the domain boundary as a closed triangular mesh. SGLDBench utilizes MATLAB’s voxelization capabilities [37] with a user-defined voxel resolution to compute a solid voxelization. For complex-shaped simulation domains, the voxels are classified as solid or void, depending on the centroid of the voxel. Alternatively, users can upload a 3D voxel grid that discretizes the domain and marks each voxel as solid or void. Void elements are excluded from the finite element analysis.
Voxels with at least one of their 26 neighboring voxels classified as void are designated as boundary voxels. SGLDBench applies a dilation operation to expand the boundary by assigning any voxel adjacent to an initial boundary voxel as a new boundary voxel. This voxelized object serves as the foundation for all subsequent operations in SGLDBench.
Boundary Conditions. The boundary conditions define where the object is fixed and how loads are applied. Fixations and forces are assigned to the nodes of the boundary elements. The user specifies the extent and position of an axis-aligned box or sphere, and SGLDBench automatically fixes or applies the specified loads to all boundary nodes within this region. Similarly, the nodes to be reset can be selected in the same manner. When all nodes of a finite element are fixed, the element becomes rigid and does not respond to any loads.
Passive Elements. Passive elements are used to preserve specific geometric features, such as object boundaries or notches for mounting connections. Passive elements remain solid throughout the optimization process and contribute to the stiffness of the structure. SGLDBench supports two general methods for specifying passive elements: Setting all boundary elements as passive, which is common in infill design problems, and setting elements with loaded or fixed nodes as passive to preserve geometric features during optimization. The dilation operation can also be used to enlarge passive structures as needed. Figure 3 illustrates the transformation from a boundary mesh to a voxel model, including boundary conditions and passive elements.
Figure 3: (a) A triangle mesh defines the domain boundary. (b) After voxelizing the domain, the user specifies boundary forces and fixes grid vertices. Black and violet dots indicate fixed vertices and vertices subjected to an external force, respectively. (c) Passive elements are shown in brown.